Increase 360 in the ratio 5:6?????????????????????????

Grade:12

3 Answers

Arun
6 years ago
Dear Shivangi
 
I didn't get your question. Please elaborate.
I will be happy to help you.
 
 
Regards
Arun (askIITians forum expert)
Priyansh
6 years ago
i didnt get your question perhaps you have missed out something..................................................................................................
jyoti bhatia
5 years ago

The three numbers are in the ratio 4:5:6.

So, let the numbers be 4x, 5x and 6x.

Now the sum is given as 360.

So,

4x +5x +6x = 360

=> 15x = 360

=> x= 24

Therefore the numbers are,

4×24 = 96

5×24 = 120

6×24 = 144